The math teacher asked Edith, " how old are you?" Sixty years less than five time my brother's age," she answered. " that doesn't help me," replied the teacher. " yes, it does," said Edith, " he and i are twins!" How old is Edith?
Answer by swincher4391(1107)

5x -60 = x
-60 = -4x
x =15